# From Rankings to Insights: Evaluation Should Shift Focus from Leaderboard to Feedback

### v2

@@ B.1 Experimental Setup for Evaluation Method
Prompt of Naive Criteria Decomposition Pointwise Evaluation You are an impartial judge responsible for evaluating the quality of responses provided by different LLMs to a given [question]. Your task is to design a comprehensive evaluation framework that includes clearly defined metrics and their respective weights. You shuold answer step by step. You should answer in English. Please carefully follow these steps:
1. **Analyze Question**: You must first analyze the question. The objective of this comparison is to find important factors that significantly influence the quality of the responses.
2. **Develop Metrics**: Establish a hierarchical set of evaluation metrics. There should be 3 to 9 primary metrics. Each primary metric should have several detailed sub-metrics to provide specific, measurable criteria for evaluating the responses.
3. **Assign Weights**: Allocate appropriate weights to each metric based on its relative importance in distinguishing the quality of the responses. The weights should be integers, and the sum of all weights should equal 100.
4. **Output Format**: Present the final evaluation framework in a structured list format. You do not need to include the primary metrics; only the secondary metrics are required, in the following format:
<Evaluation_Framework>
1. Description of Secondary Metric 1 | Weight 1
2. Description of Secondary Metric 2 | Weight 2
3. Description of Secondary Metric 3 | Weight 3
<Evaluation_Framework>
[User Question]
{question}
@@ B.2 More Experiment on PC2-Pointwise Evaluation
In this section, we present the accuracy of PC2 pointwise evaluation on the RMB [62] benchmark. The results shown in Tab. 4 demonstrate that while PC2 maintains the pointwise evaluation paradigm, it achieves results comparable to pairwise evaluation.
Table 4: Comparison of accuracy across different LLM-as-a-Judge methods on RMB [62].
Method Overall Generation Chat Role Playing Brainstorming Summarization Rewrite Classification Translation Open QA Closed QA Code Reasoning
Pairwise Evaluation 0.7427 0.8000 0.6566 0.7949 0.7469 0.7544 0.7374 0.8333 0.7267 0.7014 0.6923 0.8280 0.7331
Vanilla Pointwise Evaluation 0.6197 0.6723 0.5060 0.7436 0.5768 0.6316 0.5354 0.7222 0.6333 0.6082 0.5692 0.7640 0.5657
PC2 Pointwise Evaluation 0.7353 0.8000 0.6687 0.8333 0.7676 0.6667 0.6667 0.7778 0.6867 0.7315 0.6718 0.7600 0.7689
   + Baseline Answer 0.7394 0.7940 0.7108 0.7821 0.7083 0.6842 0.6263 0.6667 0.6846 0.7890 0.7077 0.7800 0.7490
@@ B.3 Experimental Setup for Evaluation and Leaderboard
Models. We select 17 mainstream LLMs for evaluation. For closed-source models, we select claude3.7-sonnet-20250219 [65], gpt-4o-2024-11-20 [28], qwen-max-2024-10-15 , doubao-1-5-pro-32k-250115 [35], and hunyuan-standard-2025-02-10 [66]. For open-source models, we select deepseek-v3-250324 [34], deepseek-r1-250120 [36], QwQ-32B [31], Gemma-3-27B-IT [67], Qwen2.5-72B-Instruct [68], Gemma-3-4B-IT [67], Qwen2.5-32B-Instruct [68], Meta-Llama-3.1-70B-Instruct [69], Qwen2.5-7B-Instruct [68], Meta-Llama-3.1-8B-Instruct [69], Mistral-7B-Instruct-v0.3 , and Phi-4-Mini-Instruct [70]. Among these, deepseek-v3-250324 and deepseek-r1-250120 are deployed with fp8, while the other models are deployed with bf16. For all models, if their official decoding parameters are provided, we use the default settings. If not, we set the temperature to 0.7 and top_p to 0.95.
@@ B.4 Leaderboard
We propose in the paper how to quickly check the weaknesses of a LLM through the tree structure. Due to the limitation of the paper’s length, we only present the first-level nodes (i.e., the six domain) in the paper, while the more fine-grained results are presented on our project website. The leaderboard results, as presented in Tab. 5, reveal some valuable weaknesses of existing LLMs. These weaknesses can significantly guide the model’s development if they are timely available during the development process of the model. Deepseek-v3-250324 ranks at the forefront among all tested models due to its superior coding, mathematics, and reasoning capabilities. However, it is noteworthy that its roleplay ability exhibits significant deficiencies. Claude3.7-sonnet-20250219 demonstrates high proficiency in coding, yet its overall capabilities are relatively weak. Gemma-3-4B-IT achieves exceptionally high scores in liberal arts subjects, but it performs poorly in coding and mathematics, which are more science-oriented. Additionally, it is worth mentioning that doubao-1-5-pro-32k-250115 possesses a notable advantage in mathematical abilities.
@@
Model Size Overall Writing Roleplay Knowledge Coding Mathematics Reasoning
Closed-source and Open-source (>100B) LLMs
deepseek-v3-250324 671B 454.47(1) 454.08(3) 443.75(5) 464.54(2) 452.89(1) 450.28(1) 464.53(1)
deepseek-r1-250120 671B 453.76(2) 468.82(1) 461.19(1) 460.32(4) 434.31(3) 418.86(3) 456.35(4)
[…]
Open-source (<100B) LLMs
QwQ-32B 32B 452.75(3) 465.50(2) 455.78(2) 464.94(1) 426.15(4) 430.91(2) 462.08(3)
Gemma-3-27B-IT 27B 437.61(4) 449.74(4) 452.91(3) 462.22(3) 399.08(6) 392.74(4) 462.13(2)
Qwen2.5-72B-Instruct 72B 395.04(7) 406.01(8) 378.90(9) 409.28(8) 383.07(7) 372.91(8) 407.08(9)
Gemma-3-4B-IT 4B 391.92(9) 420.14(7) 431.55(6) 426.82(7) 326.42(12) 292.43(14) 421.03(7)
Qwen2.5-32B-Instruct 32B 370.52(11) 386.67(12) 356.79(10) 380.65(11) 342.44(10) 360.14(10) 388.93(11)
Meta-Llama-3.1-70B-Instruct 70B 362.23(12) 388.10(11) 360.94(9) 368.04(12) 336.79(11) 305.13(13) 381.05(12)
Qwen2.5-7B-Instruct 7B 357.59(13) 378.48(13) 348.52(12) 367.58(13) 324.79(13) 339.03(11) 369.78(13)
Meta-Llama-3.1-8B-Instruct 8B 338.07(15) 366.84(14) 345.72(13) 345.94(15) 310.33(15) 271.33(16) 341.63(16)
Mistral-7B-Instruct-v0.3 7B 314.22(16) 351.71(16) 331.22(15) 324.97(16) 272.39(16) 223.40(17) 324.49(17)
Phi-4-Mini-Instruct 3.8B 309.55(17) 328.02(17) 320.75(16) 307.78(17) 268.48(17) 287.93(15) 345.47(15)
@@ C.1.2 Initial Annotation for Query Type Tags
Initial low-quality Type Labeling ### Task
Your task is to categorize a given query based on the following rules. For each rule, explain your reasoning before assigning a label. Evaluate the query against each domain in the order provided below until a match is found. Once a domain is identified, proceed with its specific sub-rules. If no domain matches, proceed to the next domain check. If no specific domain applies, return a generic "other" label.
#### Rules
##### Step 1: Determine the Domain
Evaluate the query against the following domains in this order. Stop and proceed with the corresponding sub-rules once a domain is matched. If the domain is labeled "other", return {"domain": "other", "domain_name": "<custom_domain_name>"} and do not proceed further. Explain your reasoning before assigning a label.
1. Writing - Condition: If the query relates to writing, purposeful professional writing, literature, storytelling, grammar, or language-related topics. - Label: "writing" - If not matched, proceed to the next domain.
2. Roleplay - Condition: If the query involves roleplay, character interactions, storytelling, or immersive scenario-based dialogue. - Label: "roleplay" - If not matched, proceed to the next domain. 3. Coding - Condition: If the query relates to programming, algorithms, or code-related topics. - Label: "coding" - If not matched, proceed to the next domain. 4. Mathematics - Condition: If the query pertains to mathematical concepts, computations, proofs, or problem-solving. - Label: "mathematics" - If not matched, proceed to the next domain. 5. Reasoning - Condition: If the query pertains to reasoning, logic, critical thinking, or problem-solving without direct reference to specific knowledge or programming. - Label: "reasoning" - If not matched, proceed to the next domain. 6. Knowledge - Condition: If the query pertains to factual knowledge or general subject areas such as science, history, literature, philosophy, or current affairs. - Label: "knowledge" - If not matched, label as "other" with a custom "domain_name" and stop. #### Step 2: Domain-Specific Categorization
Once the domain is determined, apply the corresponding sub-rules below. If the domain is "other", skip this step. Explain your reasoning before assigning a label.
- Domain: "writing" - Label 2: Writing Topic - Specify the writing topic in lowercase (e.g., "creative writing", "speech", "LLM prompt", "game", "academic", "social media", "shopping", "formal", "casual", etc.). - If unclear or not explicitly mentioned, use "null". - If none of the predefined topics fit, provide a custom description. - Label 3: Task Type - Choose applicable task types (multiple may apply): - "Creative Writing": Crafting original content, poems, or creative pieces. - "Polishing": Improving clarity, style, and flow. - "Grammar Check": Identifying and correcting grammar mistakes. - "Content Refining": Enhancing word choice, tone, or style. - "Summarization": Condensing text while retaining the core message. - "Expanding": Adding details or elaboration. - "Paraphrasing": Restating content differently. - "Writing Feedback": Providing constructive feedback. - "Translating": Converting text between languages. - "Tone Adjustment": Modifying tone for audience/context. - "Text Summarization": Analyzing meaning, themes, or key points. - If none fit, provide a custom description.
@@
- Domain: "roleplay" - Label 2: Genre - Specify the genre in lowercase (e.g., "sci-fi", "fantasy", "historical", "romance", "mystery", "horror", "action/adventure", "supernatural", etc.). - If multiple genres apply, list them all. - If none fit, provide a custom description. - Label 3: Task Type - Choose applicable task types (multiple may apply): - "Dialogue Generation": Creating natural character conversations. - "Content Creation": Writing from a character’s perspective. - "Question Answering": Responding within character constraints. - "Emotional Simulation": Modeling character-appropriate emotions. - "Decision Support": Providing character-based guidance. - "Skill Training": Simulating expert/practice partners. - "Behavior Simulation": Replicating character behaviors. - "Evaluation & Feedback": Assessing from a character’s perspective. - If none fit, provide a custom description. - Domain: "coding" - Label 2: Programming Language - Specify the language in lowercase (e.g., "python", "java", "c++", "sql", etc.). - If unclear or not mentioned, use "null". - Label 3: Task Type - Choose applicable task types (multiple may apply): - "Code Generation": Creating code from a prompt. - "Debugging": Resolving code errors. - "Code Explanation": Explaining code functionality. - "Code Editing": Modifying existing code. - "Conceptual QA": Answering programming theory questions. - "Code Translation": Converting code between languages. - "Bug Fixing": Fixing specific code issues. - "Testing": Writing/analyzing tests. - "Data Structures & Algorithms": Solving related problems. - "System Design": Designing software architecture. - "Optimization": Improving code efficiency. - "Security": Addressing vulnerabilities. - "Documentation": Writing/improving documentation. - If none fit, provide a custom description. - Domain: "mathematics" - Label 2: Subfield - Specify the subfield in lowercase (e.g., "algebra", "calculus", "geometry", "statistics & probability", etc.). - If unclear, use "null". - Label 3: Task Type - Choose applicable task types (multiple may apply): - "Computation": Performing calculations. - "Proof Writing": Writing formal proofs. - "Conceptual Understanding": Explaining concepts. - "Problem-Solving": Applying techniques to challenges. - "Real-World Application": Connecting math to scenarios. - "Graphing & Visualization": Interpreting/drawing graphs. - "Logical Deduction": Deriving conclusions logically. - If none fit, provide a custom description. - Domain: "reasoning" - Label 2: Reasoning Type - Specify the type in lowercase (e.g., "logical reasoning", "critical thinking", "problem-solving", "analytical thinking", etc.). - If unclear, use "null". - Label 3: Task Type - Choose applicable task types (multiple may apply): - "Logical Deduction": Deriving conclusions logically. - "Puzzle Solving": Solving riddles or brainteasers. - "Argument Evaluation": Analyzing argument validity. - "Scenario Analysis": Considering scenario outcomes. - "Theoretical Reasoning": Exploring abstract concepts. - "Creative Problem-Solving": Finding novel solutions. - "Ethical Reasoning": Evaluating based on ethics. - If none fit, provide a custom description. - Domain: "knowledge" - Label 2: Subject - Specify the subject in lowercase (e.g., "history", "philosophy", "geography", "medicine", "economics", etc.). - If unclear, use "null". - Label 3: Task Type - Choose applicable task types (multiple may apply): - "Fact Recall": Retrieving specific facts. - "Conceptual Understanding": Explaining/interpreting ideas. - "Comparative Analysis": Comparing concepts. - "Causal Reasoning": Examining cause-and-effect. - "Logical Deduction": Applying logic to conclude. - "Hypothetical Thinking": Exploring "what if" scenarios. - "Problem-Solving": Solving practical/theoretical problems. - "Opinion-Based": Providing subjective reasoning. - If none fit, provide a custom description. #### Step 3: Final Output
Return a Python dictionary based on the identified domain and labels:
- For "other": {"domain": "other", "domain_name": "<custom_domain_name>"}
@@
 - Replace "task" with the list of identified task types (or empty list [] if none apply).
 - Replace <content> with the corresponding value (e.g., "poetry", "sci-fi", "history", etc.).

### v3

@@ B.1 Experimental Setup for Evaluation Method
Prompt of Naive Criteria Decomposition Pointwise Evaluation You are an impartial judge responsible for evaluating the quality of responses provided by different LLMs to a given [question]. Your task is to design a comprehensive evaluation framework that includes clearly defined metrics and their respective weights. You shuold answer step by step. You should answer in English. Please carefully follow these steps: 1. **Analyze Question**: You must first analyze the question. The objective of this comparison is to find important factors that significantly influence the quality of the responses. 2. **Develop Metrics**: Establish a hierarchical set of evaluation metrics. There should be 3 to 9 primary metrics. Each primary metric should have several detailed sub-metrics to provide specific, measurable criteria for evaluating the responses. 3. **Assign Weights**: Allocate appropriate weights to each metric based on its relative importance in distinguishing the quality of the responses. The weights should be integers, and the sum of all weights should equal 100. 4. **Output Format**: Present the final evaluation framework in a structured list format. You do not need to include the primary metrics; only the secondary metrics are required, in the following format: <Evaluation_Framework> 1. Description of Secondary Metric 1 | Weight 1 2. Description of Secondary Metric 2 | Weight 2 3. Description of Secondary Metric 3 | Weight 3 <Evaluation_Framework> [User Question] {question}
@@ B.2 More Experiment on PC2-Pointwise Evaluation
More Benchmarks. To further validate the effectiveness of our method, we conduct experiments on additional benchmarks, namely RewardBench-v2 Malik et al. (2025) and RMB Zhou et al. (2025). Both benchmarks adopt the Best-of-N evaluation paradigm, which requires judge model selecting the best answer among multiple candidates, making the task inherently more difficult. The experimental results are reported in Tab. 4. As can be seen, our method consistently improves performance across all benchmarks, thereby validating its broad effectiveness.
naive ours
RMB 0.4068 0.5969
RewardBench-v2 0.5551 0.5869
Table 4: Evaluation accuracy results on RewardBench-v2 and RMB.
Ablation Study. To gain deeper insights into the factors that contribute to the performance of our method, we further conduct extensive ablation studies along three dimensions: (i) pre-com model combination, (ii) number of pre-com models, and (iii) order of pre-com models. The results are summarized in Tab. 5, 6, and 7, respectively.
(1) Ablation on model combination. For model combination (Tab. 5), we observe that multi-model setups consistently outperform the single-model baseline (gpt-4o), confirming that a single model is insufficient for robust judging performance. However, the differences among various multi-model combinations are relatively small, indicating that the diversity of the incorporated models, rather than the different combinations of models, is the key factor that drives performance improvements.
(2) Ablation on number of models. As shown in Tab. 6, increasing the number of models does not always lead to better performance. The best result is obtained when using $3$ models, whereas adding more models sometimes introduces redundancy or noise, resulting in a slight drop in accuracy.
(3) Ablation on order of models. Finally, we investigate the effect of model ordering (Tab. 7). The results show that different ordering permutations lead to only marginal differences in accuracy, suggesting that the impact of model order is relatively limited. This indicates that our approach is robust to changes in execution sequence, and the overall performance is primarily determined by the chosen set of models rather than their specific order.
These findings provide practical guidance for constructing and optimizing multi-model judging systems, highlighting the importance of not only choosing the right set of models but also carefully designing their arrangement.
Model Combination ACC
single model (gpt-4o) 0.5974
gpt-4o + doubao-pro-1.5-32k + Deepseek-V3 0.6962
gpt-4o + doubao-pro-1.5-32k + Qwen2.5-7B-Instruct 0.6904
gpt-4o + doubao-pro-1.5-32k + Meta-Llama-3.1-8B-Instruct 0.6747
gpt-4o + doubao-pro-1.5-32k + Phi-4-mini-instruct 0.7077
gpt-4o + Qwen2.5-7B-Instruct + Meta-Llama-3.1-8B-Instruct 0.6825
gpt-4o + Qwen2.5-7B-Instruct + Phi-4-mini-instruct 0.6574
Qwen2.5-7B-Instruct + Phi-4-mini-instruct + Meta-Llama-3.1-8B-Instruct 0.6699
Table 5: Ablation on model combination.
Number of Models Model Combination ACC
2 gpt-4o + doubao-pro-1.5-32k 0.6699
3 gpt-4o + Deepseek-V3 + doubao-pro-1.5-32k 0.6962
4 gpt-4o + doubao-pro-1.5-32k + Deepseek-V3 + Qwen2.5-7B-Instruct 0.6621
5 gpt-4o + doubao-pro-1.5-32k + Deepseek-V3 + Qwen2.5-7B-Instruct + Phi-4-mini-instruct 0.6448
Table 6: Ablation on number of models.
Order of Models ACC
gpt-4o → doubao-pro-1.5-32k → Deepseek-V3 0.6668
gpt-4o → Deepseek-V3 → doubao-pro-1.5-32k 0.6962
Deepseek-V3 → gpt-4o → doubao-pro-1.5-32k 0.6699
Deepseek-V3 → doubao-pro-1.5-32k → gpt-4o 0.6920
doubao-pro-1.5-32k → gpt-4o → Deepseek-V3 0.6778
doubao-pro-1.5-32k → Deepseek-V3 → gpt-4o 0.6605
Table 7: Ablation on order of models.
@@ B.3 Experimental Setup for Evaluation and Leaderboard
Models. We select 21 mainstream LLMs for evaluation. For closed-source models, we select claude3.7-sonnet-20250219 Anthropic (2025), gpt-4o-2024-11-20 OpenAI (2024a), qwen-max-2024-10-15 , doubao-1-5-pro-32k-250115 ByteDance (2024), and hunyuan-standard-2025-02-10 Tencent (2025). For open-source models, we select Deepseek-V3-250324 DeepSeek-AI (2024), Deepseek-R1-250120 DeepSeek-AI (2025), QwQ-32B Team (2025b), Gemma-3-27B-IT Team et al. (2025), Qwen2.5-72B-Instruct Yang et al. (2024b), Gemma-3-4B-IT Team et al. (2025), Qwen2.5-32B-Instruct Yang et al. (2024b), Meta-Llama-3.1-70B-Instruct Grattafiori et al. (2024), Qwen2.5-7B-Instruct Yang et al. (2024b), Meta-Llama-3.1-8B-Instruct Grattafiori et al. (2024), Mistral-7B-Instruct-v0.3 , and Phi-4-Mini-Instruct Abdin et al. (2024). Additionally, we evaluate recent open-source models: GPT-OSS-20B OpenAI (2025), GPT-OSS-120B OpenAI (2025), Qwen3-8B Team (2025a), and Qwen3-32B Team (2025a). Among these, Deepseek-V3-250324 and Deepseek-R1-250120 are deployed with fp8, while the other models are deployed with bf16. For all models, if their official decoding parameters are provided, we use the default settings. If not, we set the temperature to 0.7 and top_p to 0.95.
@@ B.4 Leaderboard
We propose in the paper how to quickly check the weaknesses of a LLM through the tree structure. Due to the limitation of the paper’s length, we only present the first-level nodes (i.e., the six domain) in the paper, while the more fine-grained results are presented on our project website. The leaderboard results, as presented in Tab. 8, reveal some valuable weaknesses of existing LLMs. These weaknesses can significantly guide the model’s development if they are timely available during the development process of the model.
The new overall leader, GPT-OSS-120B, exhibits a strong aptitude for technical tasks, securing top ranks in coding, reasoning, and mathematics. However, its performance is less dominant in creative domains like roleplay. A clear trade-off is visible in other models; for example, Deepseek-R1-250120 excels more in writing and roleplay but is weaker in mathematics. Conversely, Gemma-3-27B-IT achieves the top score in knowledge and ranks highly in reasoning, yet its performance is notably lower in STEM-oriented subjects like coding and mathematics. Despite a lower overall ranking, claude-3-7-sonnet-20250219 shows competitive coding abilities, but its weaker reasoning skills hinder its general performance. Additionally, it is worth mentioning that doubao-1-5-pro-32k-250115 possesses a notable advantage in mathematical abilities relative to its overall rank.
Model Overall Writing Roleplay Knowledge Coding Mathematics Reasoning
Closed-source and Open-source (>100B) LLMs
GPT-OSS-120B 242.24(1) 233.59(4) 231.33(7) 247.99(3) 263.35(1) 241.59(2) 235.70(1)
Deepseek-V3-250324 237.65(3) 233.04(5) 237.41(5) 250.91(2) 247.09(3) 220.01(5) 229.93(4)
Deepseek-R1-250120 235.25(4) 244.14(1) 248.29(2) 238.19(4) 231.94(6) 190.58(8) 234.91(2)
claude-3-7-sonnet-20250219 225.07(7) 232.74(6) 224.94(10) 229.79(8) 238.37(5) 187.36(10) 203.56(11)
gpt-4o-2024-11-20 214.80(9) 225.90(8) 234.42(6) 219.31(10) 202.03(10) 171.82(11) 210.50(9)
qwen-max-2024-10-15 182.59(13) 187.44(13) 174.14(14) 192.56(13) 183.33(13) 164.28(13) 178.68(14)
doubao-1-5-pro-32k-250115 182.23(14) 175.36(18) 173.19(15) 192.40(14) 187.05(12) 192.35(6) 181.46(13)
hunyuan-standard-2025-02-10 156.44(19) 163.21(20) 144.43(20) 156.85(18) 158.49(18) 149.47(15) 152.47(19)
Open-source (<100B) LLMs
Qwen3-32B 241.13(2) 241.29(2) 246.03(3) 234.47(6) 246.87(4) 245.29(1) 224.93(6)
Gemma-3-27B-IT 229.88(5) 236.63(3) 243.59(4) 253.25(1) 207.01(9) 192.19(7) 233.82(3)
QwQ-32B 229.38(6) 232.26(7) 250.41(1) 237.71(5) 224.73(7) 187.67(9) 228.24(5)
Qwen3-8B 222.88(8) 224.63(9) 227.69(9) 210.07(11) 219.59(8) 241.40(3) 216.47(7)
GPT-OSS-20B 213.72(10) 196.23(12) 176.09(13) 230.79(7) 251.59(2) 225.26(4) 212.92(8)
Gemma-3-4B-IT 203.43(11) 224.30(10) 230.57(8) 219.62(9) 167.27(15) 137.90(17) 207.15(10)
Qwen2.5-72B-Instruct 190.42(12) 197.99(11) 179.91(11) 194.18(12) 190.53(11) 171.61(12) 194.16(12)
Qwen2.5-32B-Instruct 176.10(15) 185.19(14) 170.58(16) 174.20(15) 174.39(14) 162.31(14) 173.53(15)
Meta-Llama-3.1-70B-Instruct 170.51(16) 183.13(15) 178.49(12) 167.73(16) 165.25(16) 133.37(18) 170.07(16)
Qwen2.5-7B-Instruct 166.67(17) 179.10(16) 164.64(17) 162.77(17) 161.99(17) 146.67(16) 163.15(17)
Meta-Llama-3.1-8B-Instruct 159.01(18) 176.65(17) 159.26(18) 152.30(20) 154.93(19) 119.67(20) 159.35(18)
Mistral-7B-Instruct-v0.3 148.94(20) 163.44(19) 158.07(19) 152.55(19) 138.24(21) 114.74(21) 133.61(21)
Phi-4-mini-instruct 143.99(21) 150.48(21) 144.18(21) 142.36(21) 143.21(20) 123.93(19) 148.08(20)
@@ C.1.2 Initial Annotation for Query Type Tags
Initial low-quality Type Labeling ### Task Your task is to categorize a given query based on the following rules. For each rule, explain your reasoning before assigning a label. Evaluate the query against each domain in the order provided below until a match is found. Once a domain is identified, proceed with its specific sub-rules. If no domain matches, proceed to the next domain check. If no specific domain applies, return a generic "other" label. #### Rules ##### Step 1: Determine the Domain Evaluate the query against the following domains in this order. Stop and proceed with the corresponding sub-rules once a domain is matched. If the domain is labeled "other", return {"domain": "other", "domain_name": "<custom_domain_name>"} and do not proceed further. Explain your reasoning before assigning a label. 1. Writing - Condition: If the query relates to writing, purposeful professional writing, literature, storytelling, grammar, or language-related topics. - Label: "writing" - If not matched, proceed to the next domain. 2. Roleplay - Condition: If the query involves roleplay, character interactions, storytelling, or immersive scenario-based dialogue. - Label: "roleplay" - If not matched, proceed to the next domain. 3. Coding - Condition: If the query relates to programming, algorithms, or code-related topics. - Label: "coding" - If not matched, proceed to the next domain. 4. Mathematics - Condition: If the query pertains to mathematical concepts, computations, proofs, or problem-solving. - Label: "mathematics" - If not matched, proceed to the next domain. 5. Reasoning - Condition: If the query pertains to reasoning, logic, critical thinking, or problem-solving without direct reference to specific knowledge or programming. - Label: "reasoning" - If not matched, proceed to the next domain. 6. Knowledge - Condition: If the query pertains to factual knowledge or general subject areas such as science, history, literature, philosophy, or current affairs. - Label: "knowledge" - If not matched, label as "other" with a custom "domain_name" and stop. #### Step 2: Domain-Specific Categorization Once the domain is determined, apply the corresponding sub-rules below. If the domain is "other", skip this step. Explain your reasoning before assigning a label. - Domain: "writing" - Label 2: Writing Topic - Specify the writing topic in lowercase (e.g., "creative writing", "speech", "LLM prompt", "game", "academic", "social media", "shopping", "formal", "casual", etc.). - If unclear or not explicitly mentioned, use "null". - If none of the predefined topics fit, provide a custom description. - Label 3: Task Type - Choose applicable task types (multiple may apply): - "Creative Writing": Crafting original content, poems, or creative pieces. - "Polishing": Improving clarity, style, and flow. - "Grammar Check": Identifying and correcting grammar mistakes. - "Content Refining": Enhancing word choice, tone, or style. - "Summarization": Condensing text while retaining the core message. - "Expanding": Adding details or elaboration. - "Paraphrasing": Restating content differently. - "Writing Feedback": Providing constructive feedback. - "Translating": Converting text between languages. - "Tone Adjustment": Modifying tone for audience/context. - "Text Summarization": Analyzing meaning, themes, or key points. - If none fit, provide a custom description. - Domain: "roleplay" - Label 2: Genre - Specify the genre in lowercase (e.g., "sci-fi", "fantasy", "historical", "romance", "mystery", "horror", "action/adventure", "supernatural", etc.). - If multiple genres apply, list them all. - If none fit, provide a custom description. - Label 3: Task Type - Choose applicable task types (multiple may apply): - "Dialogue Generation": Creating natural character conversations. - "Content Creation": Writing from a character’s perspective. - "Question Answering": Responding within character constraints. - "Emotional Simulation": Modeling character-appropriate emotions. - "Decision Support": Providing character-based guidance. - "Skill Training": Simulating expert/practice partners. - "Behavior Simulation": Replicating character behaviors. - "Evaluation & Feedback": Assessing from a character’s perspective. - If none fit, provide a custom description. - Domain: "coding" - Label 2: Programming Language - Specify the language in lowercase (e.g., "python", "java", "c++", "sql", etc.). - If unclear or not mentioned, use "null". - Label 3: Task Type - Choose applicable task types (multiple may apply): - "Code Generation": Creating code from a prompt. - "Debugging": Resolving code errors. - "Code Explanation": Explaining code functionality. - "Code Editing": Modifying existing code. - "Conceptual QA": Answering programming theory questions. - "Code Translation": Converting code between languages. - "Bug Fixing": Fixing specific code issues. - "Testing": Writing/analyzing tests. - "Data Structures & Algorithms": Solving related problems. - "System Design": Designing software architecture. - "Optimization": Improving code efficiency. - "Security": Addressing vulnerabilities. - "Documentation": Writing/improving documentation. - If none fit, provide a custom description. - Domain: "mathematics" - Label 2: Subfield - Specify the subfield in lowercase (e.g., "algebra", "calculus", "geometry", "statistics & probability", etc.). - If unclear, use "null". - Label 3: Task Type - Choose applicable task types (multiple may apply): - "Computation": Performing calculations. - "Proof Writing": Writing formal proofs. - "Conceptual Understanding": Explaining concepts. - "Problem-Solving": Applying techniques to challenges. - "Real-World Application": Connecting math to scenarios. - "Graphing & Visualization": Interpreting/drawing graphs. - "Logical Deduction": Deriving conclusions logically. - If none fit, provide a custom description. - Domain: "reasoning" - Label 2: Reasoning Type - Specify the type in lowercase (e.g., "logical reasoning", "critical thinking", "problem-solving", "analytical thinking", etc.). - If unclear, use "null". - Label 3: Task Type - Choose applicable task types (multiple may apply): - "Logical Deduction": Deriving conclusions logically. - "Puzzle Solving": Solving riddles or brainteasers. - "Argument Evaluation": Analyzing argument validity. - "Scenario Analysis": Considering scenario outcomes. - "Theoretical Reasoning": Exploring abstract concepts. - "Creative Problem-Solving": Finding novel solutions. - "Ethical Reasoning": Evaluating based on ethics. - If none fit, provide a custom description. - Domain: "knowledge" - Label 2: Subject - Specify the subject in lowercase (e.g., "history", "philosophy", "geography", "medicine", "economics", etc.). - If unclear, use "null". - Label 3: Task Type - Choose applicable task types (multiple may apply): - "Fact Recall": Retrieving specific facts. - "Conceptual Understanding": Explaining/interpreting ideas. - "Comparative Analysis": Comparing concepts. - "Causal Reasoning": Examining cause-and-effect. - "Logical Deduction": Applying logic to conclude. - "Hypothetical Thinking": Exploring "what if" scenarios. - "Problem-Solving": Solving practical/theoretical problems. - "Opinion-Based": Providing subjective reasoning. - If none fit, provide a custom description. #### Step 3: Final Output Return a Python dictionary based on the identified domain and labels: - For "other": {"domain": "other", "domain_name": "<custom_domain_name>"}
@@
- Replace "task" with the list of identified task types (or empty list [] if none apply). - Replace <content> with the corresponding value (e.g., "poetry", "sci-fi", "history", etc.).